OpenStudy (anonymous):

Can someone check my calc prob and tell me if my answer is correct? Find the derivative of the function. y=Ine^x

OpenStudy (anonymous):

y(prime)=1/x(e^x)+e^x(In) and my answer is y(prime)=e^x(1/x+In)

OpenStudy (anonymous):

y=lne^x=e^x de^x/dx =e^x

OpenStudy (anonymous):

u lost me

OpenStudy (anonymous):

hold the phone \(\ln(e^x)=x\)

OpenStudy (anonymous):

well the problem has absolutely No parenthesis

OpenStudy (anonymous):

sorry do u knw dat log n exp are mirror fns and that why they cancel out each other

OpenStudy (anonymous):

ya is it 1

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

is the correct ans 1

OpenStudy (anonymous):

i have no idea this is an even problem and my book does not carry evens numbered problems

OpenStudy (unklerhaukus):

\[y=\ln_ex\] ?

OpenStudy (anonymous):

No simply lne^x

OpenStudy (anonymous):

lne^x = xlne = x So y = x Now derivative it.

OpenStudy (anonymous):

y=In e^x

OpenStudy (anonymous):

see ur que was to diff y= lne^x now as i said ln and exp cancels out each other so lne^x=x ie y=x dx/dx=1

OpenStudy (anonymous):

you guys are completely confusing me

OpenStudy (anonymous):

did u get tis

OpenStudy (anonymous):

Do you know logarithm properties?

OpenStudy (anonymous):

one question, is it possible to cancel well when rewritten.....ln=1/w so we have 1/x e^x is it possible to cancel out the x's?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

i mean ln=1/x

OpenStudy (anonymous):

\[\ln a^b = b \ln a \]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

So \[\ln e^x = x \ln e\]

OpenStudy (anonymous):

And ln e = 1

OpenStudy (anonymous):

Do you know logarithm properties?

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

you can do that way too but rem tis log n exponential r mirrors function dats y can cancel them out and wat micahwood50 says is also correct

OpenStudy (anonymous):

Well, so y = x. Now derivative it.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yeah. y' = 1

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

No problem. Glad I helped you.

OpenStudy (anonymous):

:]